1.Find the equation of the normal to the parabola y^2+4x=0 at the point where the line y=x+c touches it. Q.(2) Identify and trace the conic x^2-2xy+y^2-3x+2y+3=0.
1
Expert's answer
2019-04-09T08:57:47-0400

A)Firstly we will find a normal vector to the

"y-x-c = 0;"

n(-1,1). Let’s find intersection parabola and line.


"y^2 + 4y - 4c = 0;"


Line touches parabola, the discriminant is equal 0


"D = 16 - 16c = 0"


"c = 1"

Let A(x,y) - intersection of line and parabola.


"y = x -1 \lor y^2 +4x = 0"

A(-1,-2)

So the equation of line is

"X = A+k*n;"

- vector-equation of line.

B)The maximum power of equation is 2. So it is conic anyway. 


"x^2-2xy+y^2-3x+2y+3=0."


"Ax^2 + Bxy + Cy^2+Dx + Ey + F = 0"

eccentricity is equal

"\sqrt{2*\sqrt{(a-c)^2+b^2}/((a+c)^2+\sqrt{(a-c)^2+b^2})}"

Eccentricity is equal 2^0.5 and it is greater than 1. So it is hyperbola